OpenStudy (anonymous):

Find x for -sin 2x = 0?

OpenStudy (arthur_ser):

-sin2x=0 -sin2x=sin0 -2x = 2nπ + 0 or -2x= 2nπ + π - 0 , where n is an integer simplify that